### v2

# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

Abstract.
Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems. Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques. However, these techniques have several drawbacks (e.g., unsatisfactory performance). Large Language Models (LLMs) have garnered significant attention in the Software Engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks. Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks. However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear. Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for code smell detection. In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for code smell detection. To begin, we constructed high-quality Java code smell datasets sourced from GitHub. We then fine-tuned four small LMs (SLMs) and five LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection. Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory. Our analysis further explored the performance of SLMs versus LLMs in the context of code smell detection. Generally, LLMs outperform SLMs on detecting Complex Conditional, Complex Method, and Feature Envy. However, SLMs demonstrate superior performance on detecting Data Class. We also investigated the impact of varying hyper-parameter settings and the availability of training data. Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection. In contrast, we observed significant performance gains when the training dataset size was increased. Finally, our results show that LMs with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and general-purpose LLMs under In-Context Learning in the task of code smell detection. Specifically, we observe improvements in MCC ranging from 4.11% to 19.25% across the four types of code smells studied, demonstrating a more effective solution for this task. In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for code smell detection. The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for code smell detection. We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
Code Smell Detection, Large Language Model, Parameter-Efficient Fine-Tuning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
1. Introduction
Code smells are indicators of potential issues in code or design (fowler1999refactoring, ), which could adversely affect software quality, particularly in terms of maintainability (palomba2018maintainability, ), code readability (abbes2011empirical, ), and testability (chatzigeorgiou2010investigating, ). Detecting code smells can be used to identify software design issues that lead to potential problems in future development and maintenance (fowler1999refactoring, ; lanza2007object, ; moha2010decor, ). Manual identification of code smells can be time-consuming and costly (prokic2024prescriptive, ), prompting many researchers to explore automated methods for more efficient code smell detection. Existing automated code smell detection methods are typically heuristics-based (marinescu2004detection, ; morales2017on, ; moha2010decor, ). Traditional heuristics-based detectors for code smells rely on a set of manually curated features and thresholds to discriminate between smelly and non-smelly classes (azeem2019machine, ). While it has been shown that these static detectors provide reasonably accurate results, they are proven to have limitations that hinder their practical application (fernandes2016review, ; zhang2011code, ), particularly because their performance is highly dependent on the selection of thresholds. Besides, code smells identified by such heuristics-based methods can be subjectively interpreted by developers, leading to a significant gap between detection results and developers’ agreement (azeem2019machine, ; fontana2016antipattern, ; mantyla2006subjective, ; schumacher2010building, ).
To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; fontana2016comparing, ; sharma2021code, ; liu2021deep, ). By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ). However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs). The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain. Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ). However, directly using LLMs without any task-specific fine-tuning for code smell detection yields suboptimal results (mesbah2025leveraging, ; silva2024detecting, ; sadik2025benchmarking, ), as these models lack adaptation to the unique demands of the task.
Fine-tuning enables the adaptation of Language Models (LMs) to specific downstream tasks by updating model parameters, allowing them to better capture task-relevant features and improve performance. Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited. PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts. Previous studies have explored the efficiency of PEFT methods for LMs on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), etc. However, to the best of our knowledge, no prior research has focused on improving the performance of LLMs for code smell detection, nor has any in-depth study been conducted analyzing the application of PEFT methods in this area. Additionally, identifying the code smells selected in this study requires both method- and class-level code understanding. While existing research on PEFT methods has predominantly focused on code-related tasks at the method level (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), a significant gap remains in assessing their effectiveness at the class level. By leveraging the task of code smell detection, our study aims to fill this gap by systematically exploring the capacities of LMs with PEFT methods in class-level code understanding. Furthermore, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection. Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ). As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells. In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code. For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell. This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task. In this paper, we define small LMs (SLMs) as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered LLMs. Specifically, we construct high-quality Java code smell datasets comprising four types of widely studied and sophisticated code smells at both the method and class level, i.e., Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC). These datasets are developed to address the lack of suitable existing resources, as most available datasets either do not conform to the required data formats or are based on outdated codebases. We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular SLMs and five cutting-edge LLMs with parameter sizes ranging from 125M to 7B. The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage. Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the code smell detection task. Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
The main contributions of this work are as follows:
  • We mine Java repositories from GitHub and employ traditional heuristic-based detectors, along with a two-stage manual review, to create high-quality and balanced datasets for four types of code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class.
  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection. Our results indicate that LLMs, when fine-tuned, do not always outperform SLMs. Moreover, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning. Overall, LMs with PEFT methods exhibit promising capabilities in method- and class-level code understanding.
  • We explore the performance of PEFT methods under different hyper-parameter settings and in low-resource scenarios. Our results reveal that adjusting hyper-parameters to increase trainable parameters does not necessarily lead to better effectiveness in detecting code smells. However, the performance of PEFT methods improves with more training samples.
  • We compare a heuristic-based detector, two best-in-class DL-based detection tools, and two general-purpose LLMs using In-Context Learning (ICL) against our PEFT-based method on code smell detection. Our results demonstrate that LMs with PEFT methods consistently outperform all baselines across multiple code smell types, highlighting the superiority of PEFT methods.
  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on code smell detection.
The remainder of this paper is organized as follows: Section 2 presents the background information. Section 3 outlines the research questions and the process of dataset construction, while the study results are provided in Section 4. Section 5 discusses the insights for practitioners. Potential threats to validity are addressed in Section 6, with related work reviewed in Section 7. Finally, Section 8 concludes the paper and suggests future directions for research.
2. Background
2.1. The Need of Applying Language Models on Code Smell Detection
Existing heuristic-based code smell detectors rely on pre-defined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification. However, these tools often fail to consider the inherent context of the code, which is crucial for accurately detecting and understanding code smells (sharma2021code, ). For example, the Java method shown in Figure 1 contains a switch-case statement with seven cases. From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week, from Monday to Sunday. It is straight-forward and easy to understand, as all the case statements follow the same structure. However, when scanned by an existing heuristic-based code smell detector, such as DesigniteJava (designitejava, ), the tool flags the method with the Complex Method code smell. This decision is based solely on the method’s cyclomatic complexity, which exceeds a pre-defined threshold, without considering the method’s context.
In contrast, Language Models (LMs), such as ChatGPT (achiam2023gpt, ), have the advantage of considering the inherent context of the code (wang2021codet5, ; roziere2024codellama, ). The contextual understanding allows the LM to accurately determine that the method does not exhibit Complex Method, as it can perform a role similar to a skilled developer assessing the code. The LM can evaluate not only traditional metrics, such as cyclomatic complexity, but also analyze additional factors, including readability, modularity, and overall design. Therefore, compared to traditional heuristic-based static analysis tools and certain ML- or DL-based detectors, LMs have the potential to detect code smells more accurately by considering both structural and contextual factors, rather than relying solely on numerical thresholds or pre-defined rules.
Figure 1. An example of a Java method that is considered to contain Complex Method code smell by the heuristic-based tool
2.2. Full Fine-tuning
Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre, ). Full fine-tuning involves updating all parameters of an LM, including multi-head attention (MHA), feed-forward neural network (FFN), layer normalization, and so on. However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey, ; ding2023parameter, ).
2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ).
The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. These methods were selected based on several considerations. First, they are all supported by the widely used Hugging Face PEFT library [^1] https://github.com/huggingface/peft, which provides a standardized and efficient framework for applying PEFT to LMs. Second, according to the PEFT library’s documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$. The four selected methods span these categories, making them representative of the major approaches currently available. Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring, ; li2024comprehensive, ; mannisto2025comparative, ; zhou2024empirical, ; taylor2022galactica, ; esmaeili2024empirical, ; wistuba2024choice, ). Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study. By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
Prompt tuning (lester2021power, ) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer. When fine-tuning PLMs with prompt tuning, only the soft prompts are updated, while the remaining model weights and architecture are kept fixed (not updated).
Prefix tuning (li2021prefix, ) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers. Specifically, these trainable prefix vectors are prepended to the key $K$ and value $V$ of all the MHA mechanisms. Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter, ).
LoRA (Low-Rank Adaptation) (hu2022lora, ) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA. For a pre-trained weight matrix $W_{0}\in\{W_{q},W_{k},W_{v}\}$, an additive operation $W_{0}+\Delta W=W_{0}+BA$ is performed, where the rank of matrices $B$ and $A$ is considerably smaller than that of $W_{0}$, consequently effectively reducing the number of trainable parameters while preserving the essential characteristics of the original matrices.
$\mathbf{(IA)^{3}}$ (liu2022few, ) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN. The operations conducted within the MHA can be outlined as follows:
[…]
where $\odot$ denotes the element-wise multiplication and $d_{k}$ refers to the dimension of key $K$.
The modifications in the FFN can be defined as:
[…]
where $\sigma$ is the activation function within FFN, and $W_{up}$ and $W_{down}$ indicate the weight matrices of the FFN.
3. Methodology
Figure 2. Left: architecture of our proposed method for code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
3.1. Research Goal
Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods. We frame the code smell detection task as a binary classification problem, where the input is a piece of code. A sequence classification head is added on top of the LLM to handle the classification task. For model fine-tuning, we use four distinct PEFT methods. The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
3.2. Research Questions
This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in SLMs and LLMs for code smell detection. To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
RQ3: What is the effectiveness of PEFT methods in low-resource scenarios? Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks. Due to the scarcity of real-world datasets (zakeri2023systematic, ), these scenarios are common in tasks such as code smell detection. Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristic-based and DL-based methods. Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large, ; shin2023prompt, ; li2025large, ; xu2024unilog, ), and such models can also be leveraged for code smell detection. With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
3.3. Dataset Construction
In this section, we elaborate on the process with the following steps of constructing the high-quality and balanced datasets for four types of code smells used to fine-tune both small and large LMs for code smell detection. We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ). While some code smell datasets provide valid code samples, they do not cover the specific types of code smells we focus on (e.g., Complex Conditional and Complex Method) (madeyski2020mlcq, ). To better align with current software development practices, we constructed new datasets based on trending GitHub projects created within the past six years. Furthermore, since data imbalance may negatively impact model performance on code smell detection (khleel2024improving, ; nandini2025improving, ; fontana2016comparing, ; nunes2024tuning, ), we ensured that our constructed datasets contain an equal number of positive and negative samples. Figure 3 illustrates the overview of the dataset construction process.
Figure 3. Overview of dataset construction
3.3.1. Selected Code Smells
In this study, we intend to use LMs with PEFT methods to detect four types of code smells, comprising three at the method level (i.e., Complex Conditional, Complex Method, and Feature Envy) and one at the class level (i.e., Data Class). The definitions of these four code smell types are provided below:
We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, ; fontana2016comparing, ; barbez2020machine, ; kim2017finding, ; liu2023deep, ; liu2018deep, ; skipina2024automatic, ; yang2024feature, ; fontana2013code, ; madeyski2020mlcq, )). Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring, ). For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring, ; sjoberg2012questioning, ; khomh2009exploratory, ).
Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava, ) and PMD (pmd, ), which facilitates the collection of candidate code samples potentially exhibiting these smells. We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics. In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code, ), which expands the ambition and enhances the applicability of this study.
At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding. CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell. However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring, ; tsantalis2009indentification, ; liu2023deep, ; alkharabsheh2021analysing, ). Therefore, it also demands class-level code understanding. The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
3.3.2. Java Repository Selection
To construct a code smell dataset that accurately represents modern coding practices, we selected Java repositories created in the past six years from GitHub. We used the GitHub REST API [^3] https://docs.github.com/en/rest?apiVersion=2022-11-28 to gather available Java projects that meet the following criteria:
  • The repositories should be created between January 1, 2019 and April 30, 2024.
  • The repositories should exclude forked ones to avoid duplicates.
  • The repositories should have a minimum of 1,000 stars and 1,000 lines of code.
These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial. By employing these stringent filtering criteria, we obtained a list of 401 GitHub repositories coded in Java.
3.3.3. Potential Code Smell Detection
DesigniteJava is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD is a rule-based static analysis tool that identifies common programming issues, including code smells. Both of them have been extensively used in prior studies (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ; fernandes2016review, ; aniche2018code, ; wu2024ismell, ). In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC. The output of this step is a list of detected code smells, each annotated with the corresponding project, package, class, and, where applicable, method name. It is worth noting that the code smell detection results from DesigniteJava and PMD may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting certain type of code smell, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
3.3.4. Data Preparation
CodeSplitJava (codesplitjava, ) is a utility program designed to extract methods or classes from Java source code and organize them into separate files. We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.3.2, generating individual code files for each method or class. The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders). By aligning the list of detected code smells obtained in Section 3.3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristic-based detectors as either containing code smells or being free of them.
Table 1. Number of potential positive code samples after each step of data preparation (#Initial Samples shows the initially collected potential positive samples, #After Deduplication reflects the number of samples after duplicate removal, and #After Token Limit indicates the count of samples that pass the token length filter, i.e., no more than 1,024 tokens)
Code Smell #Initial Samples #After Deduplication #After Token Limit
Complex Conditional 23,984 7,768 6,901
Complex Method 20,171 14,687 9,819
Feature Envy 2,152 1,113 337
Data Class 9,672 8,501 7,121
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. This decision was made to accommodate the input constraints of the models used in our experiments (see Section 3.6). Specifically, the input window size for the LMs was set to 1,024 tokens to align with GPU memory limitations. Samples exceeding 1,024 tokens undergo truncation, which may result in the loss of critical contextual information. For instance, key sections of code relevant to the detection of code smells could be omitted, potentially impacting the model’s performance. Therefore, we set a hard limit of 1,024 tokens to avoid such issues, ensuring that all samples in our dataset can be fully utilized by the models without sacrificing the integrity of their content. Table 1 provides the number of potential positive samples after each step for the four types of code smells. Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample. This decision aligns with the definition of FE (see Section 3.3.1), which emphasizes the method’s interaction with its enclosing class. Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
3.3.5. Manual Review
After the above steps, we obtained two sets of samples for each code smell: a potential positive set and a potential negative set, with the latter being significantly larger in size. To construct a high-quality and balanced dataset, the first and the fourth authors carried out a two-stage manual review process, as illustrated in Figure 4.
In the first stage, we reviewed all potential positive samples manually. Each sample was carefully examined to identify and eliminate false positives. If a sample in the potential positive set had been incorrectly labeled as positive by DesigniteJava or PMD, it was reclassified as negative and moved to the negative set; otherwise, it was transferred into the positive set. This stage yielded a verified positive set and a smaller verified negative set. In the second stage, in order to enlarge the negative set to achieve dataset balance, we randomly selected samples from the potential negative set for manual validation. Similar to the first stage, samples confirmed as true negatives were added to the negative set, while those identified as false negatives were reclassified into the positive set. This process was repeated iteratively until an equal number of manually reviewed positive and negative samples was achieved. During this stage, all randomly selected potential negative samples were required to be unique, adhere to the predefined maximum token limit, and be representative. To ensure the representativeness of negative samples across the four types of code smells, we established the following criteria: for CC, negative samples must contain at least one conditional statement; for CM, negative samples must contain at least 15 lines of code; and for FE and DC, negative samples must contain at least 30 lines of code, as we extract the corresponding class for these two smells. These constraints help filter out trivial or uninformative code snippets and ensure that the negative samples are sufficiently complex and realistic, thereby contributing to a more robust and reliable dataset. Through this two-stage review, we ensured that the final dataset for each code smell type was both balanced and of high quality.
Figure 4. The process of two-stage manual review in dataset construction
Table 2. Number of positive and negative samples after manual review
Code Smell #Positive #Negative Total
Complex Conditional 4,604 4,604 9,208
Complex Method 5,748 5,748 11,496
Feature Envy 307 307 614
Data Class 6,922 6,922 13,844
In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process. For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristic-based tools were randomly selected. The two authors independently examined these samples and then compared their labeling results. Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to assess inter-rater agreement between the two authors. The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87. All values exceeded the threshold of 0.7, indicating a strong level of agreement between the annotators. Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
  • CC: Labeled as positive if the method contains at least one conditional statement that either (i) has a nesting depth of three or more, making the method hard to read, or (ii) includes compound boolean expressions involving three or more operands, exhibiting complex logic that makes the condition challenging to evaluate and understand.
  • CM: Labeled as positive if the method simultaneously satisfies both (i) three or more execution paths with complex control flow, and (ii) an implementation that is difficult to comprehend or follow.
  • FE: Labeled as positive if the method inside the input class satisfies both (i) over 50% of its data access operations, such as method calls and field accesses, target members of external classes rather than its own class, and (ii) relocating the method to one of these external classes would improve encapsulation and cohesion.
  • DC: Labeled as positive if the class contains (i) primarily fields accompanied by trivial getters/setters (at least 80% of its fields), and (ii) lacks methods that implement significant logic or domain behavior, indicating its role as a passive data holder.
These criteria were grounded in established definitions from prior studies (sharma2017house, ; fowler1999refactoring, ; fontana2016comparing, ; pietrzak2006leveraging, ) and refined based on insights from our pilot study. During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total. Whenever questions or ambiguities arose, the second author, who is an expert in code smell detection, was consulted. In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency. The final count of positive and negative samples, following the manual review, is presented in Table 2.
3.3.6. Data Split
In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ). For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset. Table 3 outlines the number of code samples for each code smell type after the data split. We then used these clean and high-quality datasets as our final datasets for a series of experiments.
Table 3. Number of positive and negative samples after data split
Code Smell Train Valid Test Total
Complex Conditional #Positive 3,684 460 460 4,604
#Negative 3,684 460 460 4,604
Complex Method #Positive 4,600 574 574 5,748
#Negative 4,600 574 574 5,748
Feature Envy #Positive 245 31 31 307
#Negative 245 31 31 307
Data Class #Positive 5,538 692 692 6,922
#Negative 5,538 692 692 6,922
3.4. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods for the task of code smell detection. To ensure a thorough analysis, we included both small and large LMs as our base models. Consistent with the definition adopted by Weyssow et al. (weyssow2023exploring, ), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
Table 4. Selected LMs and their corresponding parameter size and architecture
Model #Params Architecture
SLMs CodeBERT 125M Encoder-only
GraphCodeBERT 125M
UnixCoder 126M
CodeT5 220M Encoder-decoder
LLMs StarCoderBase-1B 1.1B Decoder-only
StarCoderBase-3B 3B
DeepSeek-Coder-1.3B 1.3B
DeepSeek-Coder-6.7B 6.7B
CodeLlama-7B 7B
First, the selected LMs should be open-source as PEFT methods require access to their parameters. Second, the models should support classification tasks as we frame code smell detection as a binary classification problem. Third, the base models used in this study must be able to be fine-tuned and tested on a single 40GB GPU (our maximum computational resource) without encountering memory overflow. Fourth, the selected models should either be popular SLMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 5 LLMs (StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments. To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only. Although decoder-only models were originally designed for generative tasks (vaswani2017attention, ; radford2019language, ), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers, ; kariyappa2024progressive, ). Therefore, we also included decoder-only models as base models in our study. Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ). Table 4 presents the details of the 9 base models, including their parameter sizes and architectural types.
3.5. Metrics
We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing, ). These metrics are calculated using the following formulas:
(3) $\text{Precision}=\frac{TP}{TP+FP}$
(4) $\text{Recall}=\frac{TP}{TP+FN}$
[…]
(6) $\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
where $TP$, $TN$, $FP$, $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
3.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024. The choice to limit the input length to 1,024 tokens was driven by hardware constraints. As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient, ), this setting enables all selected base models to train within available GPU memory capacity. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
Table 5 shows the hyperparameter settings of different PEFT methods. We adopted a batch size of 1 for all experiments to avoid GPU out-of-memory issues. We used a fixed random seed (42) across all runs to ensure experimental reproducibility. For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training. This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix, ; weyssow2023exploring, ). As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving, ). For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers. In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3). We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details can be found.
Table 5. Hyperparameter settings of different PEFT methods and full fine-tuning
PEFT Method Hyperparameter Setting Batch Size Seed
Prompt Tuning Number of virtual tokens: 20 1 42
Prefix Tuning
LoRA Rank $r$: 8
$alpha$: 16
$dropout$: 0.1
Target modules: $Q$, $K$, and $V$
$(IA)^{3}$ Target modules: $Q$, $K$, $V$, and FFN
Full Fine-tuing None
4. Results
In this section, we first illustrate the experiment setup for each RQ formulated in Section 3, and we then present the results and findings for the four RQs.
4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
4.1.1. Experiment Setup
Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes. Each PEFT method was employed to fine-tune the 9 base models (see Section 3.4). Note that we only applied LoRA and $(IA)^{3}$ to Code-T5 as the current PEFT library does not support fine-tuning encoder-decoder models with prompt tuning and prefix tuning. We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 40GB of GPU memory. Table 6 and Table 7 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC, CM, FE, and DC, respectively.
Table 6. Results of small LMs and LLMs with different PEFT methods on detecting CC and CM (bold: highest values per model, underline: highest values among all models)
Model Method CC CM
F1 MCC F1 MCC
SLMs CodeBERT Full FT 70.54% 44.41% 80.57% 61.19%
Prompt Tuning 62.38% 25.90% 76.13% 54.00%
Prefix Tuning 63.88% 27.87% 81.33% 62.94%
LoRA 73.04% 46.10% 71.00% 49.79%
$(IA)^{3}$ 69.44% 40.49% 81.32% 62.96%
GraphCodeBERT Full FT 74.01% 48.08% 80.66% 61.33%
Prompt Tuning 65.24% 30.81% 78.90% 58.60%
Prefix Tuning 69.82% 39.97% 79.18% 60.46%
LoRA 73.15% 46.30% 80.03% 60.93%
$(IA)^{3}$ 72.50% 45.00% 80.00% 60.45%
UnixCoder Full FT 73.04% 46.10% 82.39% 64.89%
Prompt Tuning 69.35% 38.70% 78.66% 57.95%
Prefix Tuning 71.63% 43.26% 82.71% 65.84%
LoRA 72.19% 46.04% 73.18% 51.52%
$(IA)^{3}$ 73.47% 46.98% 82.23% 64.47%
CodeT5 Full FT 73.01% 46.19% 81.13% 62.74%
LoRA 33.33% 0.00% 83.29% -2.95%
$(IA)^{3}$ 43.38% -4.05% 40.32% 16.70%
LLMs StarCoderBase-1B Prompt Tuning 71.07% 42.21% 80.05% 60.11%
Prefix Tuning 70.10% 40.23% 76.88% 54.01%
LoRA 41.41% -1.62% 45.41% 7.01%
$(IA)^{3}$ 74.27% 49.01% 80.90% 62.01%
StarCoderBase-3B Prompt Tuning 69.12% 38.29% 81.87% 63.85%
Prefix Tuning 67.81% 35.69% 78.25% 56.92%
LoRA 40.13% 7.93% 68.41% 46.00%
$(IA)^{3}$ 74.68% 49.96% 82.07% 64.81%
Deepseek-Coder-1.3B Prompt Tuning 70.21% 40.44% 81.77% 63.75%
Prefix Tuning 69.38% 39.62% 79.06% 58.36%
LoRA 69.44% 40.49% 78.42% 57.96%
$(IA)^{3}$ 72.50% 45.01% 81.87% 63.82%
Deepseek-Coder-6.7B Prompt Tuning 67.93% 35.87% 81.19% 63.11%
Prefix Tuning 68.80% 37.62% 80.09% 60.57%
LoRA 69.98% 40.05% 81.34% 63.63%
$(IA)^{3}$ 70.84% 42.36% 83.34% 66.88%
CodeLlama-7B Prompt Tuning 68.68% 37.92% 81.25% 63.43%
Prefix Tuning 53.41% 15.95% 80.77% 62.08%
LoRA 70.12% 40.70% 80.70% 62.59%
$(IA)^{3}$ 70.48% 42.32% 83.00% 66.17%
Table 7. Results of small LMs and LLMs with different PEFT methods on detecting FE and DC (bold: highest values per model, underline: highest values among all models)
Model Method FE DC
F1 MCC F1 MCC
SLMs CodeBERT Full FT 74.09% 48.80% 81.98% 65.60%
Prompt Tuning 56.17% 13.07% 81.56% 64.57%
Prefix Tuning 62.82% 25.93% 81.81% 65.43%
LoRA 67.44% 36.17% 83.17% 68.48%
$(IA)^{3}$ 66.49% 38.48% 83.30% 68.18%
GraphCodeBERT Full FT 77.40% 54.95% 82.66% 67.42%
Prompt Tuning 66.88% 37.49% 83.75% 68.95%
Prefix Tuning 74.19% 48.39% 89.26% 79.26%
LoRA 72.23% 46.36% 85.85% 72.40%
$(IA)^{3}$ 69.84% 45.48% 85.58% 72.56%
UnixCoder Full FT 77.04% 56.76% 83.51% 69.49%
Prompt Tuning 77.32% 55.30% 83.68% 68.74%
Prefix Tuning 75.80% 51.64% 83.89% 69.28%
LoRA 70.97% 41.94% 83.60% 68.65%
$(IA)^{3}$ 73.95% 49.32% 83.96% 69.44%
CodeT5 Full FT 66.49% 38.48% 83.46% 69.12%
LoRA 54.07% 10.02% 70.03% 49.06%
$(IA)^{3}$ 54.65% 9.76% 46.79% -2.66%
LLMs StarCoderBase-1B Prompt Tuning 72.23% 46.36% 84.66% 71.21%
Prefix Tuning 75.65% 52.30% 85.44% 72.19%
LoRA 78.58% 60.68% 79.30% 58.85%
$(IA)^{3}$ 73.95% 49.32% 84.96% 71.71%
StarCoderBase-3B Prompt Tuning 68.96% 39.74% 83.63% 68.47%
Prefix Tuning 67.61% 35.78% 86.56% 74.06%
LoRA 70.85% 42.29% 83.70% 68.64%
$(IA)^{3}$ 69.35% 38.73% 84.90% 71.46%
DeepSeek-Coder-1.3B Prompt Tuning 78.98% 58.34% 84.17% 69.99%
Prefix Tuning 66.88% 37.49% 82.44% 64.90%
LoRA 78.98% 58.34% 82.72% 67.68%
$(IA)^{3}$ 83.87% 67.74% 85.84% 72.57%
DeepSeek-Coder-6.7B Prompt Tuning 78.89% 58.83% 85.56% 72.72%
Prefix Tuning 80.56% 61.81% 83.03% 68.14%
LoRA 77.32% 55.30% 83.43% 68.64%
$(IA)^{3}$ 77.32% 55.30% 85.34% 72.34%
CodeLlama-7B Prompt Tuning 72.52% 45.37% 85.49% 72.55%
Prefix Tuning 63.57% 30.67% 82.30% 64.61%
LoRA 77.32% 55.30% 83.10% 68.31%
$(IA)^{3}$ 74.19% 48.39% 83.54% 68.45%
4.1.2. SLMs
Overall, GraphCodeBERT proved the most effective in identifying CC and DC, while UnixCoder excelled at detecting CM and FE. In comparison, CodeT5 fine-tuned with PEFT methods performed the worst across all types of code smells. CodeBERT performed similarly to GraphCodeBERT, albeit slightly worse. For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively. This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert, ). Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding. When fine-tuned with PEFT methods, the MCC values of CodeBERT even became negative, reaching -4.05% for LoRA on CC and -2.66% for $(IA)^{3}$ on DC. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection, highlighting the limitations of its PEFT adaptation. In contrast, the fully fine-tuned CodeT5 demonstrated acceptable performance in code smell detection. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
4.1.3. LLMs
DeepSeek-Coder-1.3B stands out as the top-performing LLM, consistently achieving the highest or on-par F1-scores and MCC values across all fine-tuning methods when compared to other LLMs. Although the StarCoderBase models perform similarly to the DeepSeek-Coder models in detecting FE and DC, which involve class-level code understanding, their results are notably less stable for detecting CC and CM, which require method-level understanding. In particular, the performance of StarCoderBase models fine-tuned with LoRA degrades significantly on the detection of CC and CM. Notably, when fine-tuned with PEFT techniques, models with larger parameter sizes within the same family do not surpass, and in some cases, even slightly underperform their smaller counterparts. For example, in FE detection, StarCoderBase-1B and DeepSeek-Coder-1.3B generally achieve higher F1-scores and MCC values than StarCoderBase-3B and DeepSeek-Coder-6.7B.
4.1.4. SLMs vs. LLMs
For the detection of CC and CM, SLMs and LLMs demonstrate generally comparable performance. Notably, in the case of DC, GraphCodeBERT, a representative SLM, significantly outperforms all other models, including LLMs. These results suggest that smaller models can remain competitive, or even superior, for certain types of code smells. In contrast, for FE, LLMs fine-tuned with PEFT methods substantially outperform SLMs, likely due to the more complex and semantic nature of this smell type, which benefits from the richer contextual representations and broader knowledge captured by larger models.
Figure 5. Peak GPU memory usage for detecting CC: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
Additionally, Figure 5 summarizes the peak GPU memory usage of both SLMs and LLMs with different PEFT methods on CC detection. Comparing Figure 5 (a) and (b), we can observe that PEFT methods impose substantially higher peak GPU memory requirements on LLMs compared to SLMs, particularly for larger-sized LLMs.
4.1.5. Method-level vs. Class-level Code Understanding
The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.3.1). By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding. Notably, the models even achieve better performance on class-level code smells, as evidenced by the fact that the best-performing model yields significantly lower results on CC compared to the other smell types.
4.1.6. Best PEFT Method
Generally, for SLMs, both $(IA)^{3}$ and prefix tuning demonstrate strong performance across different models. For LLMs, $(IA)^{3}$ consistently delivers the best results on the detection of all four types of code smells. To be specific, the most effective PEFT method for SLMs on the task of code smell detection varies depending on the model: for instance, CodeBERT performs optimally with $(IA)^{3}$, whereas UnixCoder tends to achieve better results with prefix tuning. However, for LLMs, $(IA)^{3}$ always outperforms other PEFT methods across all studied code smells. Whether for StarCoderBase Models, Deepseek-Coder Models, or CodeLlama-7B, they all achieve superior performance when fine-tuned using $(IA)^{3}$. Additionally, for LLMs, prompt tuning can be employed as an alternative option. Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases. An example of this is using LoRA on StarCoderBase-1B for CM detection. A case in point is the use of LoRA on StarCoderBase-1B for detecting CM. While StarCoderBase-1B achieves a minimum MCC of 54.01% with the other three PEFT methods (i.e., prompt tuning, prefix tuning, and $(IA)^{3}$), its performance drops sharply to 7.01% with LoRA, resulting in a disparity of over 40%.
4.1.7. PEFT methods vs. Full Fine-tuning
According to Table 6 and Table 7, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection. For small LMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning for most code smell types. However, for FE, all PEFT methods perform significantly worse than full fine-tuning. These results suggest that, in most cases, PEFT methods can match or surpass the performance of standard full fine-tuning for code smell detection while adjusting fewer number of trainable parameters.
PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption. Taking CC detection as an example, Figure 5 (a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs. It can be observed that juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC. Among different PEFT methods, both $(IA)^{3}$ and prefix tuning are most effective for SLMs, while $(IA)^{3}$ consistently yields the best results for LLMs. All PEFT techniques outperform or match the performance of full fine-tuning for several code smell types, while also reducing peak GPU memory usage, demonstrating their efficiency. However, for the detection of FE, full fine-tuning remains more effective.
4.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
4.2.1. Experiment Setup
Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how SLMs and LLMs perform under different hyper-parameter settings. Specifically, we used LoRA to tune a SLM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {8, 32, 64, 640, 1,280, 2,560}. At the same time, we kept other hyper-parameters of LoRA constant, with the LoRA alpha set to 16 and the LoRA dropout set to 0.1. The experiment results are presented in Figure 6.
Figure 6. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) CC; (b) CM; (c) FE; (d) DC
4.2.2. Results of RQ2
The results in Figure 6 show that for both GraphCodeBERT and DeepSeek-Coder-1.3B, the optimal $r$ value varies across different code smell types, indicating that the effectiveness of LMs on code smell detection is sensitive to the choice of LoRA configuration. Since the LoRA $r$ value directly controls the number of trainable parameters, this observation implies that different types of code smells may require different levels of model adaptability.
Furthermore, increasing the LoRA $r$ value from 8 to 2,560, which corresponds to an increase in the number of trainable parameters, does not uniformly lead to improved performance. Although a higher $r$ value increases the number of trainable parameters, which provides the model with greater capacity to learn, the MCC scores for detecting the four code smell types do not show a monotonically increasing trend. These findings indicate that expanding the number of trainable parameters alone does not guarantee better effectiveness, emphasizing the importance of careful tuning of the LoRA configuration rather than relying solely on scaling the model.
Answer to RQ2: Increasing the number of trainable parameters by adjusting hyper-parameters does not necessarily ensure effectiveness enhancement for code smell detection task. Additionally, the effectiveness of LMs in detecting different code smells is sensitive to the configuration of PEFT methods (e.g., LoRA).
4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
4.3.1. Experiment Setup
To simulate low-resource scenarios, following previous studies (wang2023one, ; liu2023empirical, ), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset. Note that each subset is balanced, containing an equal number of positive and negative samples. We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.3.6), making it inherently a low-resource task. The experiment results are shown in Table 8.
Table 8. Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC, CM, and DC in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
Model #Training Samples Method CC CM DC
GraphCodeBERT 50 Full FT 19.23% 54.61% 62.76%
Prompt Tuning -0.90% 38.19% 59.51%
Prefix Tuning 6.47% 31.64% 50.79%
LoRA 7.29% 44.99% 62.51%
$(IA)^{3}$ 11.84% 45.15% 61.40%
100 Full FT 24.67% 51.58% 68.28%
Prompt Tuning 3.91% 39.03% 61.49%
Prefix Tuning 15.32% 38.87% 54.19%
LoRA 23.14% 46.98% 68.71%
$(IA)^{3}$ 23.66% 43.99% 65.62%
250 Full FT 28.38% 55.36% 64.93%
Prompt Tuning 13.05% 42.03% 64.69%
Prefix Tuning 20.02% 47.02% 62.19%
LoRA 31.11% 51.91% 66.79%
$(IA)^{3}$ 28.49% 52.45% 64.05%
500 Full FT 34.85% 57.03% 67.64%
Prompt Tuning 18.87% 47.50% 63.41%
Prefix Tuning 25.87% 53.05% 64.63%
LoRA 36.52% 58.31% 67.83%
$(IA)^{3}$ 34.14% 56.22% 68.58%
DeepSeek-Coder-1.3B 50 Full FT 11.38% 42.46% 68.08%
Prompt Tuning 5.88% 25.99% 48.34%
Prefix Tuning -3.70% 43.16% 41.93%
LoRA -1.48% 36.74% 71.84%
$(IA)^{3}$ 1.54% 17.95% 50.29%
100 Full FT 12.27% 24.94% 74.57%
Prompt Tuning 13.05% 33.00% 52.83%
Prefix Tuning 0.68% 45.19% 52.24%
LoRA 10.96% 46.15% 74.01%
$(IA)^{3}$ 6.55% 20.21% 59.53%
250 Full FT 26.69% 48.02% 74.61%
Prompt Tuning 22.18% 38.16% 60.06%
Prefix Tuning 14.87% 45.18% 61.13%
LoRA 20.53% 55.47% 75.33%
$(IA)^{3}$ 15.28% 35.25% 65.16%
500 Full FT 29.33% 58.89% 69.70%
Prompt Tuning 22.09% 40.15% 65.99%
Prefix Tuning 17.72% 46.98% 60.89%
LoRA 32.03% 59.06% 69.57%
$(IA)^{3}$ 9.86% 43.50% 70.78%
4.3.2. Results of RQ3
According to the results reported in Table 8, it is evident that when the data is limited to 50 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting code smells drops, with CC experiencing the most pronounced performance degradation. This significant decline in performance highlights the challenges small and large LMs face in low-resource scenarios. However, when the number of training samples of CC, CM, or DC increases to 500, both models show improved performance compared to when only 50 samples are available. This improvement underscores that a larger dataset enables both small and large LMs using PEFT methods to achieve substantially better performance in code smell detection. Moreover, with 500 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for code smell detection. When the number of training samples reaches 250 or 500, both models with PEFT techniques are able to outperform their fully fine-tuned counterparts.
Additionally, in low-resource scenarios, LoRA tends to be the most effective PEFT method. As indicated by the results of RQ1 (see Section 4.1.6), for GraphCodeBERT, prefix tuning and $(IA)^{3}$ deliver superior performance, while for DeepSeek-Coder-1.3B, $(IA)^{3}$ consistently yields the best performance across different code smells. However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest MCC values through LoRA. It suggests that under different resource settings, the most effective PEFT method may vary.
Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 50 samples. However, as the number of training samples increases, the performance of PEFT methods shows significant improvement. Moreover, when the number of training samples reaches 250 or 500, several PEFT techniques are able to outperform full fine-tuning, demonstrating their potential even in moderately low-resource settings.
Table 9. Prompt 1 for GPT-4o-mini and DeepSeek-v3 in detecting CC, CM, FE, and DC
Code Smell Prompt Template*
CC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate.
### Input: [Code Sample]
### Output: [Yes/No]
CM You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Method” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity.
### Input: [Code Sample]
### Output: [Yes/No]
FE You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Feature Envy” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Feature Envy”: This code smell occurs when a method is more interested in a class other than the one it actually is in.
### Input: [Code Sample]
### Output: [Yes/No]
DC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Data Class” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Data Class”: This code smell occurs when a class contains only fields and basic methods for accessing them.
### Input: [Code Sample]
### Output: [Yes/No]
  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage, ).
4.4. RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
4.4.1. Experiment Setup
Figure 7. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC under ICL
In this RQ, we established both heuristic-based (i.e., DesigniteJava and PMD) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )) as baselines for comparison. Additionally, In-Context Learning (ICL) (tom2020language, ) has demonstrated impressive performance in various SE tasks, such as code generation (gao2024what, ; jia2024acecoder, ). To evaluate its potential in code smell detection, we explored the effectiveness of two recent and powerful general-purpose LLMs, GPT-4o-mini (hurst2024gpt4o, ) and DeepSeek-v3 (liu2024deepseek, ), under both zero-shot and few-shot settings. We selected these two models because they represent the latest advancements in commercial and open-weight general-purpose LLMs, respectively, and have demonstrated strong performance across a range of natural language and programming-related tasks (rasheed2024task, ; bruni2025benchmarking, ; liu2024deepseek, ; jahin2025unveiling, ).
Specifically, we first designed a prompt template (i.e., prompt 1) for each of the four types of code smells examined in this study, incorporating both instructional text and smell definitions, as shown in Table 9. We then used ChatGPT (ChatGPT, ) to generate two additional variants for each template (i.e., prompt 2 and prompt 3), resulting in three distinct prompts per code smell type. For the zero-shot setting, these prompts were directly provided to GPT-4o-mini and DeepSeek-v3 without any labeled examples. In the few-shot setting, example pairs were randomly sampled from the training set, where each pair consisted of one code sample with a code smell and one without, along with their corresponding labels. These examples were combined with the base prompts to form few-shot configurations containing one, two, or three example pairs. Due to resource constraints (e.g., limited API usage budget), we restricted the few-shot setting to include no more than three example pairs per prompt. As a result, each code smell type yields a total of 12 unique prompt-example configurations, derived from three prompt variants combined with four levels of contextual information (zero-shot, plus one-, two-, and three-pair few-shot settings).
This ICL setup ensures diversity in both prompt formulations and example composition, allowing for a more comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our proposed PEFT-based methods in the context of code smell detection. Figure 7 presents the performance of GPT-4o-mini and DeepSeek-v3 on code smell detection in both zero- and few-shot settings. Table 10 summarizes the final results for RQ4, where we report only the best-performing configurations for both GPT-4o-mini and DeepSeek-v3 under ICL across all prompts and example variants.
4.4.2. ICL
It can be observed from Figure 7 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini. Additionally, both models achieve better performance under few-shot settings than in the zero-shot setting. However, their performance does not consistently improve as the number of example pairs increases.
4.4.3. Comparison with State-of-the-Art Approaches
Table 10. Results of PEFT methods compared to state-of-the-art code smell detection approaches on detecting CC, CM, FE, and DC (bold: highest values per code smell)
Code Smell Method Precision Recall F1 MCC
CC DesigniteJava 64.01% 76.96% 69.89% 34.41%
DeepSmells 51.06% 100.00% 67.61% 6.63%
AE-Dense 51.13% 49.10% 50.10% 2.92%
GPT-4o-mini 71.48% 64.46% 61.29% 35.24%
DeepSeek-v3 72.38% 71.74% 71.53% 44.12%
StarCoderBase-3B with $(IA)^{3}$ 75.18% 74.78% 74.68% 49.96%
CM DesigniteJava 64.69% 98.61% 78.12% 52.58%
DeepSmells 63.64% 79.87% 70.83% 33.35%
AE-Dense 50.54% 52.05% 51.29% 3.05%
GPT-4o-mini 79.06% 67.25% 63.54% 44.78%
DeepSeek-v3 79.74% 73.78% 72.40% 53.19%
DeepSeek-Coder-6.7B with $(IA)^{3}$ 83.52% 83.36% 83.34% 66.88%
FE DesigniteJava 73.91% 54.84% 62.96% 36.73%
DeepSmells 97.10% 51.54% 67.34% 7.48%
AE-Dense 50.82% 100.00% 67.39% 12.80%
GPT-4o-mini 77.53% 77.42% 77.40% 54.95%
DeepSeek-v3 74.29% 74.19% 74.17% 48.49%
DeepSeek-Coder-1.3B with $(IA)^{3}$ 83.87% 83.87% 83.87% 67.74%
DC PMD 76.29% 89.74% 82.47% 62.83%
DeepSmells 87.72% 67.68% 76.41% 49.55%
AE-Dense 52.79% 94.36% 67.70% 16.18%
GPT-4o-mini 84.87% 82.51% 82.21% 67.35%
DeepSeek-v3 87.57% 87.57% 87.57% 75.15%
GraphCodeBERT with prefix tuning 90.95% 90.03% 89.97% 79.26%
Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection. Additionally, among conventional baseline approaches, DeepSeek-v3 achieves the most competitive overall performance. While it shows a significant performance gap compared to LMs tuned with PEFT methods for CM and FE detection (with MCC differences exceeding 10%), its performance in CC and DC detection remains relatively strong, with only a 4%-5% reduction in MCC values, demonstrating its relative strength in these specific code smell types.
Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches. This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT techniques and ICL.
Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods reveal notable performance improvements, highlighting the effectiveness of PEFT methods for code smell detection.
5. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 10, existing methods, including heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%. Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods. While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.3.1 and Section 3.3.4). According to the recent studies (silva2024detecting, ; mesbah2025leveraging, ), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding. This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions). Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical, ; wang2023one, ; liu2024delving, ), leaving their effectiveness on class-level tasks largely unexplored. Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis. Surprisingly, as discussed in Section 4.1.5, our results reveal that LLMs with PEFT methods not only overcome the anticipated class-level challenges but even outperform method-level detection in some cases. These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE. This performance advantage aligns with established neural scaling laws (kaplan2020scaling, ), which posit that model capabilities improve predictably with increased scale. However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC. Moreover, as illustrated in Figure 5, SLMs exhibit substantially lower peak GPU memory consumption compared to LLMs, making them a more resource-efficient choice. These findings highlight a critical trade-off: while LLMs tend to yield higher absolute performance, SLMs offer a competitive and cost-effective alternative, especially in settings with limited computational resources.
Considering the model used, the code smell type, and the availability of resources when selecting the appropriate PEFT method: Based on the results of RQ1 and RQ3, we can find that for code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data. From Table 6 and Table 7, we observe that although $(IA)^{3}$ generally achieves the best performance for LLMs across most code smell types, this is not universally the case. For certain models, particularly SLMs, and specific types of code smells, prefix tuning or even prompt tuning may outperform $(IA)^{3}$. As indicated in Figure 5, the relatively high peak GPU memory consumption of $(IA)^{3}$ may limit its practicality in constrained environments, where more lightweight PEFT methods could be more suitable. Moreover, in low-resource scenarios where training data is limited, LoRA emerges as the most effective choice (see results of RQ3 in Section 4.3.2). These findings emphasize the importance of selecting PEFT strategies holistically, taking into account not only the model size and available resources, including both GPU memory and training data, but also the specific demands of the code smell being detected.
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, even when the dataset size was drastically reduced, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3. This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigate the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings. As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 4%–5% lower than those attained by the top-performing PEFT-tuned LMs. This suggests that, for certain types of code smells, well-crafted prompts and a few in-context examples may enable general-purpose LLMs to produce competitive results. Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large, ; tom2020language, ; li2025large, ), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs). However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 29.25% in MCC, respectively, compared to our PEFT-based approaches. These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey. Overall, while ICL offers a lightweight and flexible alternative, especially for rapid prototyping or scenarios with limited computational resources, PEFT techniques remain the more robust and reliable solution for achieving high-accuracy code smell detection across a broad range of smell types.
Future work could explore the use of Retrieval-Augmented Generation (RAG) for code smell detection: According to Figure 7, while models generally perform better in code smell detection when provided with randomly selected examples compared to the zero-shot setting, their performance does not consistently improve as the number of example pairs increases. This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection. By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval, ) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts. By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering. As such, future research could explore whether RAG-based approaches can provide a more effective and scalable solution to PEFT methods for code smell detection.
6. Threats to Validity
Construct Validity: One primary threat to the construct validity of this study is the construction of the datasets. We first used DesigniteJava (designitejava, ) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias. However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ; tushar2016does, ; taibi2017how, ), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
Second, the manual validation process, while aimed at ensuring dataset accuracy, could introduce some subjectivity due to its reliance on human judgment. To minimize this threat, we conducted separate pilot validations for the four types of code smells before formally annotating all the data samples. During this phase, we established and refined clear identification criteria to ensure alignment between annotators. This approach not only improved labeling consistency but also reduced potential bias in the dataset.
Lastly, due to budget constraints for commercial API usage, our evaluation of general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) under ICL settings in RQ4 was limited to a maximum of three example pairs. This sample size may not fully represent the upper bounds of ICL performance. Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance. However, prior studies such as UniLog (xu2024unilog, ) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed. As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 5 LLMs. However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings.
Furthermore, we only employed four widely used and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. While these methods cover the major categories defined in the PEFT library, the exclusion of other approaches may limit the generalizability of our findings. For example, methods such as adapter tuning (houlsby2019parameter, ) could be explored in future work to provide a more comprehensive evaluation.
Moreover, our study applied PEFT methods exclusively to the detection of Java code smells, focusing on four specific types: three method-level smells (Complex Conditional, Complex Method, and Feature Envy) and one class-level smell (Data Class). While this selection provides a certain degree of diversity, it does not fully capture the wide spectrum of code smell types encountered in real-world software systems. Expanding future work to include a wider range of smell categories would enhance the comprehensiveness of the evaluation. In addition, our current work is limited to Java. Applying PEFT techniques to detect code smells in other programming languages, such as C# or Python, could yield further insights into their generalizability and adaptability across different language ecosystems.
Internal Validity: The settings of hyper-parameters for the four PEFT methods and full fine-tuning play a crucial role in the performance of the models. Variations in these settings could lead to different results, potentially affecting the reliability of our findings. To mitigate this threat, we referred to previous works and conducted pilot experiments to determine the appropriate hyper-parameter values.
Another potential threat to internal validity stems from conducting only a single experimental run per configuration due to computational resource constraints. Although we ensured consistency across all experiments by using a fixed random seed (42), this approach cannot fully account for the inherent stochasticity in the training process. To assess the impact of randomness, we conducted supplementary experiments using two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection). The results shown in Table 11 indicate that there are only small variations in the performance metrics across different random seeds. These minor differences suggest that our core findings are robust to random initialization effects, with the overall trends remaining consistent. While some performance fluctuations are observed, they do not significantly alter the conclusions drawn from the experiments. However, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable conclusions.
Table 11. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different seeds
Seed Method Precision Recall F1 MCC
0 Full FT 73.47% 72.17% 71.79% 45.62%
Prompt Tuning 64.89% 64.89% 64.89% 29.78%
Prefix Tuning 71.32% 71.30% 71.30% 42.62%
LoRA 73.17% 73.04% 73.01% 46.21%
$(IA)^{3}$ 72.44% 72.17% 72.09% 44.61%
42 Full FT 74.06% 74.02% 74.01% 48.08%
Prompt Tuning 65.49% 65.33% 65.24% 30.81%
Prefix Tuning 70.08% 69.89% 69.82% 39.97%
LoRA 73.15% 73.15% 73.15% 46.30%
$(IA)^{3}$ 72.50% 72.50% 72.50% 45.00%
999 Full FT 73.19% 72.83% 72.72% 46.02%
Prompt Tuning 64.26% 64.24% 64.22% 28.50%
Prefix Tuning 71.21% 69.57% 68.96% 40.74%
LoRA 73.07% 72.93% 72.90% 46.00%
$(IA)^{3}$ 72.23% 72.17% 72.15% 44.41%
Additionally, data leakage represents another potential threat to internal validity. First, the poor zero-shot performance of GPT-3.5-turbo in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 4.4), as we would expect stronger zero-shot effectiveness if GPT-3.5-turbo had memorized the data. Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue. Such substantial performance gains would not occur if the models had already memorized the test sets. Furthermore, Cao et al. (cao2024concerned, ) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates. This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
7. Related Work
7.1. Code Smell Detection
Existing approaches to code smell detection predominantly rely on heuristic-based methods, which utilize predefined rules to identify problematic code patterns (pmd, ; designitejava, ; campbell2013sonarqube, ; fokaefs2007jdeodorant, ; vidal2015jspirit, ; moha2010decor, ). These methods aim to flag potential smells by analyzing the source code against established standards. For instance, PMD (pmd, ), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells. These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts. Similarly, Sharma (designitejava, ) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (fontana2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
We conduct an extensive analysis of the Machine Learning Code Quality (MLCQ) dataset, focusing on how these LLMs perform when prompted to identify and classify code smells.
Recently, researchers have started to explore the use of LLMs for code smell detection. Mesbah et al. (mesbah2025leveraging, ) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells. Silva et al. (silva2024detecting, ) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells. Sadik et al. (sadik2025benchmarking, ) created a dataset for evaluating LLMs on automated code smell detection. They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance. Wu et al. (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
Most prevailing ML-based or DL-based approaches for code smell detection train models from scratch. On the other hand, existing LLM-based approaches typically rely on simple prompting strategies without any form of fine-tuning or adaptation. Compared to their works, we harness the power of LLMs using PEFT methods to detect four types of widely studied code smells (Complex Conditional, Complex Method, Feature Envy, and Data Class), aiming to enhance the models’ ability to capture the nuanced relationships between different code smell types.
7.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks. Wang et al. (wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios. Li et al. (li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
These studies focus on evaluating PEFT methods across a range of SE tasks, primarily at the method level. Compared to their works, our work specifically targets the task of code smell detection, which, to the best of our knowledge, has not been studied before. Additionally, our work further explores the capacity of LLMs on class-level code understanding, thereby addressing a significant gap in existing research.
8. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks. Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection. Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in a wider range of SE tasks, such as bug detection, code completion, and code smell refactoring. These directions will help us better understand the generalizability and practical value of PEFT methods in real-world software development settings.
Data Availability
The replication package for this work has been made available at https://github.com/MabelQi/PEFT4CSD.
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.